# Anisotropy can make a moving active fluid membrane rough or crumpled

### v1

@@ Appendix A Linear instability analysis
If we move in the $xy$-plane along paths with a constant phase, i.e., with constant $y-x$, the exponential factor $e^{i\phi}$ does not change. Consequently, the height function $h(x,y)$ does not vary along these directions. These lines, defined by $y-x=\text{constant}$, correspond to the stripes observed in real space, forming at an angle of $\pi/4$. The periodicity of the stripes is controlled by the wavenumber $k_{\text{max}}\sim\pi/a_{0}\sim\pi$, given that $a_{0}=1$ in our simulations. A representative snapshot of this pattern can be seen in Fig. 2(a) of the main text. On the other hand, for $\alpha<-1$, the maximum growth rate occurs at $\theta=\pi/4$, corresponding to $k_{x}=k_{y}$. Now using the logic used above, we find this corresponds to the formation of stripes at an angle of $43\pi/$ 4 in the $x$-$y$ plane. A representative snapshot of this case is shown in Fig. 2(b) of the main text.
Appendix B Galilean invariance
@@
$t\rightarrow t^{\prime}\equiv t,\,x_{i}\rightarrow x_{i}^{\prime}\equiv x_{i}+\lambda_{ij}u_{j}t,\,h({\bf x},t)\rightarrow h^{\prime}({\bf x}^{\prime},t^{\prime})\equiv h({\bf x},t)+u_{i}x_{i}+\frac{\lambda_{ij}}{2}u_{i}u_{j}t,$ (15)
@@
$\frac{\partial}{\partial t}\rightarrow\frac{\partial}{\partial t^{\prime}}+\lambda_{ij}u_{j}\frac{\partial}{\partial x_{i}^{\prime}},\,\frac{\partial}{\partial x_{i}}\rightarrow\frac{\partial}{\partial x_{i}^{\prime}},$
$\frac{\partial h({\bf x},t)}{\partial t}\rightarrow\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial t^{\prime}}+\lambda_{ij}u_{j}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x_{i}^{\prime}}-\frac{\lambda_{ij}}{2}u_{i}u_{j}.$ (16)
$\frac{\lambda_{ij}}{2}\frac{\partial h({\bf x},t)}{\partial x_{i}}\frac{\partial h({\bf x},t)}{\partial x_{j}}\rightarrow\frac{\lambda_{ij}}{2}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{i}}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{j}}+\frac{\lambda_{ij}}{2}u_{i}u_{j}-\lambda_{ij}u_{j}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{i}}.$ (17)
@@
Appendix C Renormalization group calculations
@@
$S=-\int_{x,t}\hat{h}D\hat{h}+\int_{x,t}\hat{h}(\partial_{t}h+\frac{\lambda_{ij}}{2}\partial_{i}h\partial_{j}h-\nu_{ij}\partial_{i}\partial_{j}h).$ (20)
C.1 Results from the linearized equation of motion
This can be obtained from the Gaussian part of the action functional (20), i.e., by setting $\lambda_{ij}=0$ in (20).
@@ C.1 Results from the linearized equation of motion
Figure 5 shows the diagrammatic representations of the various two-point functions.
C.2 One-loop corrections to the model parameters
In the (more usual) isotropic models, applications of the Wilson momentum space RG involves integrating the wavevector magnitude $q$ within a shell $\Lambda/b\leq q\leq\Lambda,\,b>1$ Hohenberg and Halperin (1977); Täuber (2014). The present model is however anisotropic, lacking any spherical symmetry. Thus, using spherical coordinates and a spherical shell for mode elimination is not convenient for the present problem. Instead, we integrate $q_{x}$ over the whole range from $\infty$- ∞ to $\infty$, and then integrate $q_{y}$ over two thin strips $\Lambda/b\leq q_{y}\leq\Lambda$ and $-\Lambda\leq q_{y}\leq-\Lambda/b$ with $b>1$; see Fig. 6. Of course, we could have chosen another integration scheme, in which $q_{y}$ is integrated over the whole range from $\infty$- ∞ to $\infty$, and then integrate $q_{x}$ over two thin strips $\Lambda/b\leq q_{x}\leq\Lambda$ and $-\Lambda\leq q_{x}\leq-\Lambda/b$. For our model, the two schemes are equivalent, connected to the fact that the different anisotropic terms of the same order scale in the same way.
@@ C.2 One-loop corrections to the model parameters
Figure 6: Wavevector-shell integration in the one-loop diagrams obtained by perturbative expansions of the action functional (20), corresponding to the anisotropic KPZ equation. Here, $q_{x}$ is integrated over the range $\infty$- ∞ to $\infty$, while $q_{y}$ is integrated over two thin strips $\Lambda/b\leq q_{y}\leq\Lambda$ and $-\Lambda\leq q_{y}\leq-\Lambda/b$.
@@
$\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(\nu q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu q_{y}^{2})^{m}}=\frac{1}{\nu^{m}}\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(q_{x}^{2}+2\alpha q_{x}q_{y}+q_{y}^{2})^{m}}$
$=\frac{1}{\nu^{m}}\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(q_{x}-a_{+})^{m}(q_{x}-a_{-})^{m}},$ (22)
@@
where $\tilde{f}(q_{x},q_{y})$ represents functions of $q_{x}$ and $q_{y}$, and we use $\alpha={\overline{\nu}}/{\nu}$. The exponent $m$ takes values $2$ or $3$. The terms $a_{+}=q_{y}(-\alpha+ip)$ and $a_{-}=q_{y}(-\alpha-ip)$ are two poles of order $m$, where $p=\sqrt{1-\alpha^{2}}$.
@@
To evaluate this integral, we first perform the integration over $q_{x}$ in the range (,)($\infty$,$\infty$)( - ∞ , ∞ ). Subsequently, we integrate over the $y$-component of momentum within the two high-$|q_{y}|$ strips, specifically in the regions $\Lambda/b\leq q_{y}\leq\Lambda$ and $-\Lambda\leq q_{y}\leq-\Lambda/b$. See Fig. 6. For computational convenience, we define $q_{x}^{\prime}=q_{x}+\alpha q_{y}$. Under this change of variables, Eq. (22) transforms into
$\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}^{\prime}}{2\pi}\frac{\tilde{f}(q_{x}^{\prime},q_{y})}{(q_{x}^{\prime 2}+p^{2}q_{y}^{2})^{m}}+\int_{-\Lambda}^{-\Lambda/b}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}^{\prime}}{2\pi}\frac{\tilde{f}(q_{x}^{\prime},q_{y})}{(q_{x}^{\prime 2}+p^{2}q_{y}^{2})^{m}}.$ (23)
@@
The first integral receives contributions from the positive values of $q_{y}$, while the second integral accounts for the negative values of $q_{y}$. Since the denominator of each of the integrals is even in $q_{x}^{\prime}$, any terms in $\tilde{f}(q_{x}^{\prime},q_{y})$ containing odd powers of $q_{x}^{\prime}$ must vanish under integration over $q_{x}^{\prime}$. Only terms with even powers of $q_{x}^{\prime}$ survive and contribute to the integral. It turns out from the explicit expressions of the various one-loop integrals that these are also even powers of $q_{y}$. Consequently, both integrals in Eq. (23) yield the same result. Hence, in the following analysis, we evaluate the integrals considering only the positive values of $q_{y}$, noting that the contributions from negative $q_{y}$ are identical.
Below we now evaluate the one-loop diagrams that account for fluctuation-induced corrections to the model parameters.
C.3 One-loop corrections to $D$
@@
$\lambda_{ij}\lambda_{mn}\int\frac{d\Omega}{2\pi}\frac{d^{2}{q}}{(2\pi)^{2}}(q_{i})(k-q)_{j}\,q_{m}\,(k-q)_{n}\frac{D^{2}}{\Big{[}\Omega^{2}+(\nu_{lk}q_{l}q_{k})^{2}\Big{]}\Big{[}(\omega-\Omega)^{2}+\big{(}\nu_{ps}(k-q)_{p}(k-q)_{s}\big{)}^{2}\Big{]}}.$ (24)
Setting $k=0$, $\omega=0$, the above integral reduces to
$\lambda_{ij}\lambda_{mn}\int_{q,\Omega}\frac{q_{i}q_{j}q_{m}q_{n}D^{2}}{\Big{[}\Omega^{2}+(\nu_{xx}q_{x}^{2}+2\nu_{xy}q_{x}q_{y}+\nu_{yy}q_{y}^{2})^{2}\Big{]}^{2}}.$ (25)
After performing the $\Omega$-integral, the contribution in (25) becomes
$\frac{\lambda_{ij}\lambda_{mn}}{4}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{D^{2}q_{i}q_{j}q_{m}q_{n}}{(\nu_{xx}q_{x}^{2}+2\nu_{xy}q_{x}q_{y}+\nu_{yy}q_{y}^{2})^{3}}.$ (26)
Using $\lambda_{xx}=\lambda_{yy}=\lambda$ and $\lambda_{xy}=\lambda_{yx}=\overline{\lambda}$ we get
$\lambda_{ij}\lambda_{mn}q_{i}q_{j}q_{m}q_{n}=\lambda^{2}q_{x}^{4}+\lambda^{2}q_{y}^{4}+4\overline{\lambda}^{2}q_{x}^{2}q_{y}^{2}+4\lambda\overline{\lambda}q_{x}^{3}q_{y}+4\lambda\overline{\lambda}q_{y}^{3}q_{x}+2\lambda^{2}q_{x}^{2}q_{y}^{2}.$
Substituting this in Eq. (26) we get the correction to $D$ as
$\frac{D^{2}}{4\nu^{3}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}}{2\pi}\Biggl{[}\frac{\lambda^{2}q_{x}^{4}+(2\lambda^{2}+4\overline{\lambda}^{2})q_{x}^{2}q_{y}^{2}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}+\frac{4\lambda\overline{\lambda}q_{x}^{3}q_{y}+4\lambda\overline{\lambda}q_{x}q_{y}^{3}+\lambda^{2}q_{y}^{4}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}\Biggl{]},$ (27)
where $a_{+}=q_{y}[-\alpha+ip]$ and $a_{-}=q_{y}[-\alpha-ip]$ are two poles of third order. We set $\nu_{xx}=\nu_{yy}=\nu$, $\nu_{xy}=\nu_{yx}=\overline{\nu}$, and $p=\sqrt{1-\alpha^{2}}$. Also defining
$\mathcal{A}=\lambda^{2},\mathcal{B}=(2\lambda^{2}+4\overline{\lambda}^{2}),\mathcal{C}=4\lambda\overline{\lambda},\mathcal{D}=4\lambda\overline{\lambda},\mathcal{E}=\lambda^{2},$
Equation (27) becomes
$\frac{D^{2}}{4\nu^{3}}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{\mathcal{A}q_{x}^{4}+\mathcal{B}q_{x}^{2}q_{y}^{2}+\mathcal{C}q_{x}^{3}q_{y}+\mathcal{D}q_{x}q_{y}^{3}+\mathcal{E}q_{y}^{4}}{(q_{x}-a_{+})^{3}(q_{x}-a_{-})^{3}}.$ (28)
Now we perform the $q_{x}$-integral first. Considering a closed contour in the upper half complex plane, the relevant poles within the closed contour are $q_{x}=a_{+}=q_{y}(-\alpha+ip)$. The residue at $q_{x}=a_{+}$ is
$\frac{1}{2!}\frac{d^{2}}{dq_{x}^{2}}\biggr{[}\frac{\mathcal{A}q_{x}^{4}+\mathcal{B}q_{x}^{2}q_{y}^{2}+\mathcal{C}q_{x}^{3}q_{y}+\mathcal{D}q_{x}q_{y}^{3}+\mathcal{E}q_{y}^{4}}{(q_{x}-a_{-})^{3}}\biggr{]}_{q_{x}=a_{+}}.$
Evaluating, this gives
$\frac{1}{2}\biggr{[}\frac{12\mathcal{A}q_{x}^{2}+2\mathcal{B}q_{y}^{2}+6\mathcal{C}q_{x}q_{y}}{(q_{x}-a_{-})^{3}}-6\frac{4\mathcal{A}q_{x}^{3}+2\mathcal{B}q_{x}q_{y}^{2}+3\mathcal{C}q_{x}^{2}q_{y}+\mathcal{D}q_{y}^{3}}{(q_{x}-a_{-})^{4}}+12\frac{\mathcal{A}q_{x}^{4}+\mathcal{B}q_{x}^{2}q_{y}^{2}+\mathcal{C}q_{x}^{3}q_{y}+\mathcal{D}q_{x}q_{y}^{3}+\mathcal{E}q_{y}^{4}}{(q_{x}-a_{-})^{5}}\biggr{]}_{q_{x}=a_{+}}.$
Substituting $q_{x}=a_{+}$ in above expression and multiplying with $2\pi i$ gives the result for the $q_{x}$-integral. Thus Eq. (28) becomes
$\frac{3D^{2}}{64\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\int_{\frac{\Lambda}{b}}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}\biggr{[}(\mathcal{A}+\mathcal{E})+\mathcal{B}\Bigl{(}\frac{1+2\alpha^{2}}{3}\Bigl{)}-(\mathcal{C}+\mathcal{D})\alpha\biggr{]}.$ (29)
After performing the $q_{y}$-integral and using the values of $\mathcal{A},\mathcal{B},\mathcal{C},\mathcal{D}$ and $\mathcal{E}$ Eq. (29) can be written as
$\frac{3D^{2}}{16\nu^{3}(1-\alpha^{2})^{\frac{5}{2}}}\frac{\ln b}{2\pi}\biggr{[}\frac{\lambda^{2}}{4}\Bigl{(}2+\frac{2}{3}+\frac{4}{3}\alpha^{2}\Bigl{)}-2\lambda\overline{\lambda}\alpha+\overline{\lambda}^{2}\Bigl{(}\frac{1+2\alpha^{2}}{3}\Bigl{)}\biggr{]}.$ (30)
Using $\alpha={\overline{\nu}}/{\nu}$, Eq. (30) becomes
$\frac{3D^{2}}{16(\nu^{2}-\overline{\nu}^{2})^{\frac{3}{2}}}\frac{\ln b}{2\pi}\biggr{[}\frac{\lambda^{2}}{4}\Bigl{(}\frac{8\nu^{2}}{3(\nu^{2}-\overline{\nu}^{2})}+\frac{4\overline{\nu}^{2}}{3(\nu^{2}-\overline{\nu}^{2})}\Bigl{)}+\overline{\lambda}^{2}\Bigl{(}\frac{\nu^{2}}{3(\nu^{2}-\overline{\nu}^{2})}+\frac{2\overline{\nu}^{2}}{3(\nu^{2}-\overline{\nu}^{2})}\Bigl{)}-\frac{2\lambda\overline{\lambda}\nu\overline{\nu}}{(\nu^{2}-\overline{\nu}^{2})}\biggr{]}.$ (31)
C.4 One-loop corrections to $\nu_{ij}$
The one-loop Feynman diagram, with a symmetry factor 8, that contributes to the fluctuation correction of $\nu$ and $\tilde{\nu}$ is shown in Fig. 8.
Figure 8: One-loop Feynman diagram that contributes to the correction of $\nu$ and $\overline{\nu}$.

### v2

@@ I Linear instability analysis
If we move in the $xy$-plane along paths with a constant phase, i.e., with constant yxy-x, the exponential factor $e^{i\phi}$ does not change. Consequently, the height function $h(x,y)$ does not vary along these directions. These lines, defined by $y-x=\text{constant}$, correspond to the stripes observed in real space, forming at an angle of $\pi/4$. The periodicity of the stripes is controlled by the wavenumber $k_{\text{max}}\sim\pi/a_{0}\sim\pi$, given that $a_{0}=1$ in our simulations. A representative snapshot of this pattern can be seen in Fig. 1(d) of the main text. On the other hand, for $\alpha<-1$, the maximum growth rate occurs at $\theta=\pi/4$, corresponding to $k_{x}=k_{y}$. Now using the logic used above, we can say this corresponds to the formation of stripes at an angle of $3\pi/4$ in the $x$-$y$ plane. We have verified this result (not shown).
II Galilean invariance
@@
$t\rightarrow t^{\prime}\equiv t,\,x_{i}\rightarrow x_{i}^{\prime}\equiv x_{i}+\lambda u_{i}t,\,h({\bf x},t)\rightarrow h^{\prime}({\bf x}^{\prime},t^{\prime})\equiv h({\bf x},t)+u_{i}x_{i}+\frac{\lambda}{2}u^{2}t,$ (18)
@@
$\frac{\partial}{\partial t}\rightarrow\frac{\partial}{\partial t^{\prime}}+\lambda u_{i}\frac{\partial}{\partial x_{i}^{\prime}},\,\frac{\partial}{\partial x_{i}}\rightarrow\frac{\partial}{\partial x_{i}^{\prime}},$
$\frac{\partial h({\bf x},t)}{\partial t}\rightarrow\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial t^{\prime}}+\lambda u_{i}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x_{i}^{\prime}}-\frac{\lambda}{2}u^{2}.$ (19)
$\frac{\lambda}{2}\frac{\partial h({\bf x},t)}{\partial x_{i}}\frac{\partial h({\bf x},t)}{\partial x_{i}}\rightarrow\frac{\lambda}{2}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{i}}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{i}}+\frac{\lambda}{2}u^{2}-\lambda u_{i}\frac{\partial h^{\prime}({\bf x^{\prime}},t^{\prime})}{\partial x^{\prime}_{i}}.$ (20)
@@
III Renormalization group calculations
@@
$S=-\int_{x,t}\hat{h}D\hat{h}+\int_{x,t}\hat{h}(\partial_{t}h+\frac{\lambda}{2}{({\bm{\nabla}}h)^{2}}-\nu_{ij}\partial_{i}\partial_{j}h).$ (23)
III.1 Results from the linearized equation of motion
This can be obtained from the Gaussian part of the action functional (23), i.e., by setting $\lambda=0$ in (23).
@@ III.1 Results from the linearized equation of motion
Fig. 4 shows the diagrammatic representations of the various two-point functions.
III.2 One-loop corrections to the model parameters
In the (more usual) isotropic models, applications of the Wilson momentum space RG involves integrating the wavevector magnitude $q$ within a shell $\Lambda/b\leq q\leq\Lambda,\,b>1$ Hohenberg and Halperin (1977); Täuber (2014). The present model is however anisotropic, lacking any spherical symmetry. Thus, using spherical coordinates and a spherical shell for mode elimination is not convenient for the present problem. Instead, we integrate $q_{x}$ over the whole range from -$\infty$ to $\infty$, and then integrate $q_{y}$ over two thin strips $\Lambda/b\leq q_{y}\leq\Lambda$ and $\Lambda\leq q_{y}\leq-\Lambda/b$ with $b>1$; see Fig. 5. Of course, we could have chosen another integration scheme, in which $q_{y}$ is integrated over the whole range from -$\infty$ to $\infty$, and then integrate $q_{x}$ over two thin strips $\Lambda/b\leq q_{x}\leq\Lambda$ and $\Lambda\leq q_{x}\leq-\Lambda/b$. For our model, the two schemes are equivalent, connected to the fact that the different anisotropic terms of the same order scale in the same way.
@@ III.2 One-loop corrections to the model parameters
Figure 5: Wavevector-shell integration in the one-loop diagrams obtained by perturbative expansions of the action functional (23), corresponding to the anisotropic KPZ equation. Here, $q_{x}$ is integrated over the range -$\infty$ to $\infty$, while $q_{y}$ is integrated over two thin strips $\Lambda/b\leq q_{y}\leq\Lambda$ and $\Lambda\leq q_{y}\leq-\Lambda/b$.
@@
$\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{m}}=\frac{1}{\nu_{x}^{m}}\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(q_{x}^{2}+2\alpha q_{x}q_{y}+\beta q_{y}^{2})^{m}}$
$=\frac{1}{\nu_{x}^{m}}\int\frac{dq_{x}}{2\pi}\int\frac{dq_{y}}{2\pi}\frac{\tilde{f}(q_{x},q_{y})}{(q_{x}-a_{+})^{m}(q_{x}-a_{-})^{m}},$ (25)
@@
where $\tilde{f}(q_{x},q_{y})$ represents functions of $q_{x}$ and $q_{y}$, and we use $\alpha={\overline{\nu}}/{\nu}$, $\beta=\nu_{y}/\nu_{x}$. The exponent $m$ takes values $2$ or $3$. The terms $a_{+}=q_{y}(-\alpha+ip)$ and $a_{-}=q_{y}(-\alpha-ip)$ are two poles of order $m$, where $p=\sqrt{\beta-\alpha^{2}}$.
@@
To evaluate this integral, we first perform the integration over $q_{x}$ in the range $(-\infty,\infty)$. Subsequently, we integrate over the $y$-component of momentum within the two high-$|q_{y}|$ strips, specifically in the regions $\Lambda/b\leq q_{y}\leq\Lambda$ and $\Lambda\leq q_{y}\leq-\Lambda/b$. See Fig. 5. For computational convenience, we define $q_{x}^{\prime}=q_{x}+\alpha q_{y}$. Under this change of variables, Eq. (25) transforms into
$\frac{1}{\nu_{x}^{m}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}^{\prime}}{2\pi}\frac{\tilde{f}(q_{x}^{\prime},q_{y})}{(q_{x}^{\prime 2}+p^{2}q_{y}^{2})^{m}}+\frac{1}{\nu_{x}^{m}}\int_{-\Lambda}^{-\Lambda/b}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}^{\prime}}{2\pi}\frac{\tilde{f}(q_{x}^{\prime},q_{y})}{(q_{x}^{\prime 2}+p^{2}q_{y}^{2})^{m}}.$ (26)
@@
The first integral receives contributions from the positive values of $q_{y}$, while the second integral accounts for the negative values of $q_{y}$. Since the denominator of each of the integrals is even in $q_{x}^{\prime}$, any terms in $\tilde{f}(q_{x}^{\prime},q_{y})$ containing odd powers of $q_{x}^{\prime}$ must vanish under integration over $q_{x}^{\prime}$. Only terms with even powers of $q_{x}^{\prime}$ survive and contribute to the integral. It turns out from the explicit expressions of the various one-loop integrals that these are also even powers of $q_{y}$. Consequently, both integrals in Eq. (26) yield the same result. Hence, in the following analysis, we evaluate the integrals considering only the positive values of $q_{y}$, noting that the contributions from negative $q_{y}$ are identical. Below, we list the integrals arising in the calculations and evaluate the one-loop diagrams that account for fluctuation-induced corrections to the model parameters. Here, $\int_{q_{x},q_{y}}\equiv\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi}\int_{-\infty}^{\infty}\frac{dq_{x}}{2\pi}$.
$I_{1}=\int_{q_{x},q_{y}}\frac{q_{x}^{2}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{2}}=\frac{1}{4\nu_{x}^{2}}\frac{\beta}{p^{3}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$
$I_{2}=\int_{q_{x},q_{y}}\frac{q_{y}^{2}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{2}}=\frac{1}{4\nu_{x}^{2}}\frac{1}{p^{3}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$
$I_{3}=\int_{q_{x},q_{y}}\frac{q_{x}q_{y}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{2}}=-\frac{1}{4\nu_{x}^{2}}\frac{\alpha}{p^{3}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$
$I_{4}=\int_{q_{x},q_{y}}\frac{q_{x}^{4}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}}=\frac{3}{16\nu_{x}^{3}}\frac{\beta^{2}}{p^{5}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$
$I_{5}=\int_{q_{x},q_{y}}\frac{q_{x}^{2}q_{y}^{2}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}}=\frac{3}{16\nu_{x}^{3}}\frac{\beta+2\alpha^{2}}{3p^{5}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$
$I_{6}=\int_{q_{x},q_{y}}\frac{q_{x}^{3}q_{y}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}}=-\frac{3}{16\nu_{x}^{3}}\frac{\alpha\beta}{p^{5}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$
$I_{7}=\int_{q_{x},q_{y}}\frac{q_{x}q_{y}^{3}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}}=-\frac{3}{16\nu_{x}^{3}}\frac{\alpha}{p^{5}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}},$
$I_{8}=\int_{q_{x},q_{y}}\frac{q_{y}^{4}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}}=\frac{3}{16\nu_{x}^{3}}\frac{1}{p^{5}}\int_{\Lambda/b}^{\Lambda}\frac{dq_{y}}{2\pi q_{y}}.$
III.3 One-loop corrections to $D$
@@
$\frac{\lambda^{2}}{4}\int\frac{d\Omega}{2\pi}\frac{d^{2}{q}}{(2\pi)^{2}}\frac{4D^{2}\times q^{4}}{\Big[\Omega^{2}+(\nu_{x}q_{x}^{2}+\nu_{y}q_{y}^{2}+2\overline{\nu}q_{x}q_{y})^{2}\Big]^{2}}.$ (27)
Where we set $k=0$, $\omega=0$. After performing the $\Omega$-integral, above integral reduces to
$\frac{\lambda^{2}D^{2}}{4}\int\frac{d^{2}q}{(2\pi)^{2}}\frac{q^{4}}{(\nu_{x}q_{x}^{2}+2\overline{\nu}q_{x}q_{y}+\nu_{y}q_{y}^{2})^{3}},$ (28)
$=\frac{\lambda^{2}D^{2}}{4}(I_{4}+2I_{5}+I_{8}),$ (29)
$=\frac{\lambda^{2}D^{2}}{4}\frac{3}{16\nu_{x}^{3}p^{5}}\frac{\ln b}{2\pi}\Big[\beta^{2}+\frac{2(\beta+2\alpha^{2})}{3}+1\Big].$ (30)
III.4 One-loop corrections to $\nu_{ij}$
The one-loop Feynman diagram, with a symmetry factor 8, that contributes to the fluctuation correction of $\nu_{x}$, $\nu_{y}$ and $\overline{\nu}$ is shown in Fig. 7.
Figure 7: One-loop Feynman diagram that contributes to the correction of $\nu_{x}$, $\nu_{y}$ and $\overline{\nu}$.